OpenStudy (pulsified333):

Use the following matrix equation instead of that found in your text to find the values of c and d which make the equation true. Use equation provided in comment (1) To make this equation true, set c= (2) And set d

OpenStudy (pulsified333):

\[c \left(\begin{matrix}17 \\4\end{matrix}\right)+d \left(\begin{matrix}4 \\ 1\end{matrix}\right)=\left(\begin{matrix}-116 \\ -27\end{matrix}\right)\]

OpenStudy (pulsified333):

@satellite73 @Directrix

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[17c+4d=-116\\ 4c+d=-27\] is what you have to solve

OpenStudy (pulsified333):

oh i didnt even think about it like that ahhahaha

OpenStudy (pulsified333):

it worked thank you @misty1212 :D